With the given parabola and slope of a line that is tangent to the parabola, find the y-int of the tangent line.
f(x)=2x^2+2x-5, tangent line has slope 1

Answer by stanbon(75887)

With the given parabola and slope of a line that is tangent to the parabola, find the y-int of the tangent line.
f(x)=2x^2+2x-5, tangent line has slope 1
------
slope at every point = f'(x) = 4x+2
If slope = 1, 4x+2 = 1 and x = -1/4
If x = -1/4, y = 2(-1/4)^2+2(-1/4)-5 = 2(1/16)-(8/16)-(80/16) = -86/16 = -4.775
-------
Find the Equation::
Form:: y = mx+b
Solve for "b" if x = -1/4, y = -4.775, and m = 1
-4.775 = 1*(-0.25) + b
----
Ans:: b = -4.525
-------------
Cheers,
Stan H.
-----------